# spacing=nonfrench Sy-FAR: Symmetry-based Fair Adversarial Robustness

### v1

@@ Abstract
Our insight is that achieving perfect parity in realistic fairness-critical tasks, such as face recognition, is often infeasible—some classes (e.g., siblings) may be highly similar, leading to more misclassifications between them. Instead, we suggest that seeking symmetry—i.e., attacks from class to would be as successful as from to —is more tractable. Intuitively, symmetry is a desirable because class resemblance is a symmetric relation in most domains. Additionally, as we prove theoretically, symmetry between individuals induces symmetry between any set of sub-groups, in contrast to other fairness notions where group-fairness is often elusive.
We develop Sy-FAR, a technique to encourage symmetry while also optimizing adversarial robustness and extensively evaluate it using five datasets, with three model architectures, including against targeted and untargeted realistic attacks. The results show Sy-FAR significantly improves fair adversarial robustness compared to state-of-the-art methods. Moreover, we find that Sy-FAR is faster and more consistent across runs. Notably, Sy-FAR also ameliorates another type of unfairness we discover in this work—target classes that adversarial examples are likely to be classified into become significantly less vulnerable after inducing symmetry.111Our code is available at: https://github.com/haneenn24/Sy-FAR-Symmetry-based-Fair-Adversarial-Robustness.
@@ 1 Introduction
Adversarial examples—perturbed inputs that lead machine-learning (ML) models to misclassify at deployment time—pose a profound challenge to ML [3, 52]. Notably, adversarial examples are not a hypothetical risk to ML, but can be implemented under real-world constraints, thus harming ML systems’ integrity in practical settings [14, 45]. For instance, adversaries may physically realize attacks against face-recognition systems via eyeglasses they can don to evade surveillance or impersonate others [45, 46].
Realizing the imminent risk to ML systems that are becoming increasingly ubiquitous, researchers have explored various directions to enhance models’ adversarial robustness (e.g., [9, 15, 21, 33, 60] ). Different classes of defenses have been proposed, ranging from inference-time countermeasures that may refine classification boundaries or derive security guarantees (e.g., [9, 21]) to adversarial training, which augments training with attacks to improve robustness (e.g., [15, 33, 54]). From these, adversarial training is particularly appealing due to significant improvements in robust accuracy (i.e., accuracy under attacks) and preservation of the time required to make inference during deployment.
@@
While defenses substantially improve adversarial robustness, they suffer from fairness issues, due to disparities between classes (e.g., individuals in the case of face recognition) or groups (e.g., genders or ethnicities) [37, 40, 59]. In particular, while defenses enhance adversarial robustness, they often do so unevenly across classes. Said differently, it is often significantly easier to produce adversarial examples for inputs from certain (source) classes than inputs from other classes. Besides ethical implications, these disparities in robust accuracy across classes—a phenomenon termed unfairness source-class adversarial robustness—have direct security implications. For example, if such disparities exist in a face-recognition system in surveillance, then certain subjects may still easily evade detection although the overall robustness seems high. As Fig. 1 shows, source-class unfairness is a significant issue in defended face-recognition models: certain subjects receive significantly lower robust accuracy than others (e.g., Antonio Banderas, first from the left, and Beyoncé, second from the right, for the adversarially trained model).
Yet another fairness issue facing defenses that, to our knowledge, we characterize for the first time (§6.3), is that of unfair target-class robustness. Specifically, we find that certain classes are more likely to be the “sinks” of misclassified adversarial examples. Put simply, adversarial examples have higher probability of being misclassified into certain classes than others. This form of unfairness directly affects security: For example, in a face-recognition system that exhibits target-class unfairness, specific individuals would be at a higher risk of impersonation than others. Hence, in addition to countering unfair source-class adversarial robustness, as has been the focus of prior work (e.g., [18, 63]), it is critical to address unfair target-class adversarial robustness.
Last, but not least, besides achieving fairness for individuals, we should also seek to enhance fairness for different groups, such as different genders or ethnicities, so no group remains disadvantaged. Prior work has demonstrated that unfair source-class robust accuracy is exhibited not only for individual classes, but also for groups [40, 37]. A core challenge arising when seeking fairness for groups is that their number may be excessively large (exponential in the number of classes), rendering it prohibitive to improve fairness across all groups [41]. Accordingly, to our knowledge, no existing technique is able to promote fair adversarial robustness for an arbitrary group that may be unknown at training time.
@@
Several methods have been proposed to ameliorate unfair unfair source-class robustness [18, 28, 32, 51, 57, 59, 63]. These roughly share the same principle: While optimizing for overall benign and robust accuracy (so these are preserved or improved compared to standard training), they also seek to directly decrease the gap in robust accuracy between classes (e.g., by assigning higher weights for classes with lower robustness) or to increase the robust accuracy of the class with minimal robustness. In doing so, leading approaches such as FAAL [63] and SpecNorm [18] substantially improve source-class fairness. However, as shown in Fig. 1, specific classes still remain highly vulnerable even under these defenses (e.g., John Travolta, fifth from the left). Notably, previous work evaluated setups where fairness and security have little implications, as they focus on artificial object classification tasks (mainly, the CIFAR datasets [22]) and impractical attacks where adversaries introduce imperceptible perturbations (bounded in -norm) to induce misclassifications. As a result, prior work’s ability to improve fairness while preserving robustness and benign accuracy (i.e., accuracy on clean inputs) in realistic fairness- and security-critical settings remains unknown. We address this gap by porting previous approaches to realistic face-recognition settings (while also evaluating them on the original setups).
@@
In this work, we argue that achieving perfect source-class fairness, as sought by prior work [18, 28, 32, 51, 57, 59, 63], may often be infeasible. For instance, assume a setting where face recognition is trained to recognize unrelated individuals as well as two siblings (or even twins). Naturally, misclassifications between the siblings would be more challenging to prevent than misclassifications between other individuals. Achieving perfect parity in this setting may even be counterproductive, as it may require reducing the robust accuracy of classes other than the siblings.
As an alternative, we propose to tackle fairness issues in adversarially robust classification via a novel technique—by encouraging symmetry between classes, as part of a framework for Symmetry-based Fair Adversarial Robustness (Sy-FAR). Besides optimizing for benign and robust accuracy, Sy-FAR trains models that would be nearly as likely to misclassify adversarial examples from class to as they are to misclassify in the other direction (§4). Intuitively, as class resemblance is typically a symmetric relation, it is natural to seek symmetry, such that misclassifications (of adversarial examples) between class pairs would be equally likely. Consequently, no class would be significantly more disadvantaged against the other. Crucially, we find that promoting symmetry boosts fair adversarial robustness from both source- and target-class perspective (§6). Furthermore, as we formally prove, ensuring symmetry for individual classes also leads to symmetry between arbitrary groups (§7). Consequently, Sy-FAR, in a way completely agnostic to group definitions, encourages symmetry between arbitrary groups and, as a by-product, enforces fair source- and target-group robustness for any arbitrary partition of classes into groups.
@@
We extensively evaluate Sy-FAR on two vision tasks (face and object recognition), using three model architectures, and two attack types (§5). Importantly, in contrast to pervious work, we consider a realistic adversary producing adversarial eyeglasses that would lead face-recognition to misclassify [45]. Thus, our work considers a real-world setting where adversarial robustness and fairness are critical. Moreover, we compare Sy-FAR empirically to FAAL [63] and SpecNorm [18], two state-of-the-art techniques.
@@
The paper proceeds as follows. §2 discusses related work, §3 lays out the threat model, and §4 presents the technical approach behind Sy-FAR. Subsequently, §5 describes our experimental setup before we proceed to presenting our results for fair (class-level) source- and target-class adversarial robustness in §6 and group fairness in §7. §8 wraps the paper with concluding remarks.
@@ 2.1 Adversarial Examples
The proliferation of ML has raised researchers’ interest in ML’s implications on system security. Researchers have shown various attacks on ML privacy (e.g., [47]), integrity (e.g., [15, 52, 4]), and availability (e.g., [48]), among others. Specifically, adversarial examples have received a special attention due to the potential risk they pose to deployed safety- and security-critical systems during deployment: Numerous attacks targeting models of varied modalities under different assumptions and with various attack objectives were proposed since the initial discovery of adversarial examples (e.g., [14, 38, 31, 45, 43, 64]).
While one class of adversarial examples leverages imperceptible adversarial perturbations, bounded in -norm, to mislead ML models [6, 12, 15, 36, 53], another class targets more realistic settings where adversaries may be more bound by practical constraints [14, 31, 43, 45, 46, 61]. For instance, in the eyeglass attack against face recognition systems [45], adversaries may need to physically realize artifacts that are robust to printing and camera noise to mislead systems under real-world constraints. Different types of attacks may have untargeted or targeted variants [39]—i.e., aiming to achieve arbitrary misclassification or misclassification into specific classes, respectively—as well as other variants suitable for other assumptions (e.g., white- and black-box [5, 29, 38]). In this work, we study the fairness properties of ML models subject to different attack types, with a particular focus on the realistic eyeglass attack.
@@
Different forms of defenses against adversarial examples have been proposed, including, but not limited to, defenses that detect attacks (e.g., [35]); filter out adversarial perturbations (e.g., [60]); smoothen the classification boundaries to reduce model vulnerability (e.g., [7, 9, 42]); verify robustness against specific adversaries (e.g., [21]); and adversarially train models to inherently increase their robustness (e.g., [25, 33]). Due to its intuitive nature, its ability to improve adversarial robustness in a practical manner against different attack types, and absence of impact on model’s inference time, adversarial training is particularly appealing and was widely studied (e.g., [15, 26, 28, 30, 33, 44, 54, 58]). Accordingly, in this work, we focus on improving the fairness properties of adversarially trained models, especially against more realistic attacks. As adversarial training is often combined with other defenses (e.g., randomized smoothing [42]), we expect that our approach and findings would also apply to defenses other than adversarial training.
@@ 2.2 Fairness and Adversarial Robustness
Algorithmic fairness, in general, and fairness in ML, specifically, have been extensively studied in recent years [55]. Researchers have defined different fairness notions between individuals and groups, ranging from precision parity to recall parity [55], studied when notions can or cannot be satisfied simultaneously (e.g., [8]), and proposed means to promote different notions of fairness (e.g., [10, 50]). In addition to ethical considerations that often motivate fairness, another key motivation behind fairness is improving the general performance of the system (e.g., benign accuracy), which sometimes stems as a side effect of improving fairness [20].
Prior works has mainly focused on studying fairness in the context of benign inputs (e.g., [1, 17, 20]). However, a line of work has studied fairness in the context of adversarial robustness—particularly, disparities in adversarial robustness between different (source) classes and groups [37, 59]. While defenses against adversarial examples such as adversarial training significantly improve robust accuracy, they do not guarantee fairness across classes or demographic groups. In high-stakes applications such as face recognition, this limitation is particularly concerning: a system may appear robust overall, yet still perform disproportionately worse on certain classes or subpopulations, exposing vulnerable groups to higher risks of misidentification or security breaches. In concurrent work, Nanda et al. [37] and Xu et al. [59] noticed that certain classes or groups may have significantly lower robust accuracy than others. These efforts highlighted the shortcomings of evaluating fairness only under benign conditions, and reinforced the need for defenses that explicitly address both robustness and fairness under attack. As noted by Rosenberg et al. [40], similar disparities are also exhibited by face-recognition models where face-obfuscation with -norm bounded perturbations to attain privacy are less effective for minority groups.
@@
Several methods have emerged to tackle fairness in adversarial settings [18, 28, 32, 51, 57, 59, 63]. In particular, Fair Robust Learning (FRL) [59] aims to improve worst-class robust accuracy through a min–max framework that reweights adversarial samples and adjusts attack strengths during trianing. Balance Adversarial Training (BAT) [51] tackles fairness by encouraging uniformity in predicted classes and equalizing attack difficulty across source classes. Fairness-Aware Adversarial Learning (FAAL) [63] improves worst-class robustness by dynamically reweighting samples and using a form of so-called Distributionally Robust Optimization (DRO) to optimize worst-case performance. Last, Confusional Spectral Regularization (SpecNorm) [18] introduces a method that penalizes the spectral norm of the confusion matrix to bound the worst-case class-level robust error. Similar to SpecNorm, Sy-FAR optimizes particular properties of the confusion matrix on adversarial examples (§4). However, Sy-FAR departs from norm-based constraints and instead focuses on directional imbalances in misclassification rates. SpecNorm and FAAL are considered the leading approaches for enhancing fair source-class adversarial robustness, thus we include them in our experiments (§5).
In contrast to prior work, Sy-FAR aims to improve fair adversarial robustness through novel means, by optimizing symmetry—while this property sometime naturally arises, misclassification patterns of adversarial examples between most class pairs are often asymmetric [34]. Moreover, in this work we identify disparities in target-class adversarial robustness as a key limitation of existing techniques and show that Sy-FAR helps make it less severe. Last, previous work has mainly focused on predefined groups (e.g., genders or races), while we show how to improve fair adversarial robustness for any arbitrary group—addressing an issue that is known to be computationally challenging for certain notions of fairness .

### v2

@@ Abstract
Our insight is that achieving perfect parity in realistic fairness-critical tasks, such as face recognition, is often infeasible—some classes (e.g., siblings) may be highly similar, leading to more misclassifications between them. Instead, we suggest that seeking symmetry—i.e., attacks from class to would be as successful as from to —is more tractable. Intuitively, symmetry is desirable because class resemblance is a symmetric relation in most domains. Additionally, as we prove theoretically, symmetry between individuals induces symmetry between any set of sub-groups, in contrast to other fairness notions where group-fairness is often elusive.
We develop Sy-FAR, a technique to encourage symmetry while also optimizing adversarial robustness and extensively evaluate it using five datasets, with three model architectures, including against targeted and untargeted realistic attacks. The results show Sy-FAR significantly improves fair adversarial robustness compared to state-of-the-art methods. Moreover, we find that Sy-FAR is faster and more consistent across runs. Notably, Sy-FAR also ameliorates another type of unfairness we discover in this work—target classes that adversarial examples are likely to be classified into become significantly less vulnerable after inducing symmetry.
@@ 1 Introduction
Adversarial examples—perturbed inputs that lead machine-learning (ML) models to misclassify at deployment time—pose a profound challenge to ML [biggio-2013, szegedy2014intriguing]. Notably, adversarial examples are not a hypothetical risk to ML, but can be implemented under real-world constraints, thus harming ML systems’ integrity in practical settings [Evtimov17Signs, sharif2016accessorize]. For instance, adversaries may physically realize attacks against face-recognition systems via eyeglasses they can don to evade surveillance or impersonate others [sharif2016accessorize, Sharif19AGNs].
Realizing the imminent risk to ML systems that are becoming increasingly ubiquitous, researchers have explored various directions to enhance models’ adversarial robustness (e.g., [Cohen19RandSmooth, goodfellow2015advtraining, Katz19Marabou, madry2018towards, Xu18Squeeze]). Different classes of defenses have been proposed, ranging from inference-time countermeasures that may refine classification boundaries or derive security guarantees (e.g., [Cohen19RandSmooth, Katz19Marabou]) to adversarial training, which augments training with attacks to improve robustness (e.g., [goodfellow2015advtraining, madry2018towards, tong2021facesec]). From these, adversarial training is particularly appealing due to significant improvements in robust accuracy (i.e., accuracy under attacks) without increasing inference-time latency.
@@
While defenses substantially improve adversarial robustness, they suffer from fairness issues, due to disparities between classes (e.g., individuals in the case of face recognition) or groups (e.g., genders or ethnicities) [nanda2021fairness, rosenberg2023fairness, xu2021robust]. In particular, while defenses enhance adversarial robustness, they often do so unevenly across classes. Said differently, it is often significantly easier to produce adversarial examples for inputs from certain (source) classes than inputs from other classes. Besides ethical implications, these disparities in robust accuracy across classes—a phenomenon termed unfair source-class adversarial robustness—have direct security implications. For example, if such disparities exist in a face-recognition system in surveillance, then certain subjects may still easily evade detection although the overall robustness seems high. As Fig. 1 shows, source-class unfairness is a significant issue in defended face-recognition models: certain subjects receive significantly lower robust accuracy than others (e.g., Antonio Banderas, first from the left, and Beyoncé, second from the right, for the adversarially trained model).
Yet another fairness issue facing defenses that, to our knowledge, we characterize for the first time (§6.3), is that of unfair target-class robustness. Specifically, we find that certain classes are more likely to be the “sinks” of misclassified adversarial examples. Put simply, adversarial examples have higher probability of being misclassified into certain classes than others. This form of unfairness directly affects security: For example, in a face-recognition system that exhibits target-class unfairness, specific individuals would be at a higher risk of impersonation than others. Hence, in addition to countering unfair source-class adversarial robustness, as has been the focus of prior work (e.g., [jin2025enhancing, zhang2024towards]), it is critical to address unfair target-class adversarial robustness.
Last, but not least, besides achieving fairness for individuals, we should also seek to enhance fairness for different groups, such as different genders or ethnicities, so no group remains disadvantaged. Prior work has demonstrated that unfair source-class robust accuracy is exhibited not only for individual classes, but also for groups [rosenberg2023fairness, nanda2021fairness]. A core challenge arising when seeking fairness for groups is that their number may be excessively large (exponential in the number of classes), rendering it prohibitive to improve fairness across all groups [rothblum2018probably]. Accordingly, to our knowledge, no existing technique is able to promote fair adversarial robustness for an arbitrary group that may be unknown at training time.
@@
Several methods have been proposed to ameliorate unfair source-class robustness [jin2025enhancing, li2023wat, ma2022fat, sun2023improving, wei2023cfa, xu2021robust, zhang2024towards]. These roughly share the same principle: While optimizing for overall benign and robust accuracy (so these are preserved or improved compared to standard training), they also seek to directly decrease the gap in robust accuracy between classes (e.g., by assigning higher weights for classes with lower robustness) or to increase the robust accuracy of the class with minimal robustness. In doing so, leading approaches such as FAAL [zhang2024towards] and SpecNorm [jin2025enhancing] substantially improve source-class fairness. However, as shown in Fig. 1, specific classes still remain highly vulnerable even under these defenses (e.g., John Travolta, fifth from the left). Notably, previous work evaluated setups where fairness and security have little implications, as they focus on artificial object classification tasks (mainly, the CIFAR datasets [krizhevsky2009learning]) and impractical attacks where adversaries introduce imperceptible perturbations (bounded in -norm) to induce misclassifications. As a result, prior work’s ability to improve fairness while preserving robustness and benign accuracy (i.e., accuracy on clean inputs) in realistic fairness- and security-critical settings remains unknown. We address this gap by porting previous approaches to realistic face-recognition settings (while also evaluating them on the original setups).
@@
In this work, we argue that achieving perfect source-class fairness, as sought by prior work [jin2025enhancing, li2023wat, ma2022fat, sun2023improving, wei2023cfa, xu2021robust, zhang2024towards], may often be infeasible. For instance, assume a setting where face recognition is trained to recognize unrelated individuals as well as two siblings (or even twins). Naturally, misclassifications between the siblings would be more challenging to prevent than misclassifications between other individuals. Achieving perfect parity in this setting may even be counterproductive, as it may require reducing the robust accuracy of classes other than the siblings.
As an alternative, we propose to tackle fairness issues in adversarially robust classification via a novel technique—by encouraging symmetry between classes, as part of a framework for Symmetry-based Fair Adversarial Robustness (Sy-FAR). Besides optimizing for benign and robust accuracy, Sy-FAR trains models that would be nearly as likely to misclassify adversarial examples from class to as they are to misclassify in the other direction (§4). Intuitively, as class resemblance is typically a symmetric relation, it is natural to seek symmetry, such that misclassifications (of adversarial examples) between class pairs would be equally likely. Consequently, no class would be significantly more disadvantaged against the other. Crucially, we find that promoting symmetry boosts fair adversarial robustness from both source- and target-class perspectives (§6). Furthermore, as we formally prove, ensuring symmetry for individual classes also leads to symmetry between arbitrary groups (§7). Consequently, Sy-FAR, in a way completely agnostic to group definitions, encourages symmetry between arbitrary groups and, as a by-product, enforces fair source- and target-group robustness for any arbitrary partition of classes into groups.
@@
We extensively evaluate Sy-FAR on two vision tasks (face and object recognition), using three model architectures, and two attack types (§5). Importantly, in contrast to pervious work, we consider a realistic adversary producing adversarial eyeglasses that would lead face-recognition to misclassify [sharif2016accessorize]. Thus, our work considers a real-world setting where adversarial robustness and fairness are critical. Moreover, we compare Sy-FAR empirically to FAAL [zhang2024towards] and SpecNorm [jin2025enhancing], two state-of-the-art techniques.
@@
The paper proceeds as follows. We next discuss related work (§2), lay out the threat model (§3), and present the technical approach behind Sy-FAR (§4). Subsequently, we describe our experimental setup (§5) before we present our results for fair (class-level) source- and target-class adversarial robustness (§6) and group fairness (§7). We wrap the paper with discussion of limitations and future work (§8) and concluding remarks (§9).
@@ 2.1 Adversarial Examples
The proliferation of ML has raised researchers’ interest in ML’s implications on system security. Researchers have shown various attacks on ML privacy (e.g., [membership-inference-paper]), integrity (e.g., [goodfellow2015advtraining, szegedy2014intriguing, Biggio12Pois]), and availability (e.g., [shumailov2021sponge]), among others. Specifically, adversarial examples have received special attention due to the potential risk they pose to deployed safety- and security-critical systems during deployment: Numerous attacks targeting models of varied modalities under different assumptions and with various attack objectives were proposed since the initial discovery of adversarial examples (e.g., [Evtimov17Signs, papernot-2016-blackbox, mahmood-binary, sharif2016accessorize, Schon18AdvSound, zouGCG-UniversalTransferableAdversarial2023]).
While one class of adversarial examples leverages imperceptible adversarial perturbations, bounded in -norm, to mislead ML models [carlini2019evaluating, Croce20Auto, goodfellow2015advtraining, Moosavi16DeepFool, szegedy-2014-intriguing], another class targets more realistic settings where adversaries may be more bound by practical constraints [Evtimov17Signs, mahmood-binary, Schon18AdvSound, sharif2016accessorize, Sharif19AGNs, Xu16PDF]. For instance, in the eyeglass attack against face recognition systems [sharif2016accessorize], adversaries may need to physically realize artifacts that are robust to printing and camera noise to mislead systems under real-world constraints. Different types of attacks may have untargeted or targeted variants [jsma-paper]—i.e., aiming to achieve arbitrary misclassification or misclassification into specific classes, respectively—as well as other variants suitable for other assumptions (e.g., white- and black-box [Brendel18DBA, liu2017delving, papernot-2016-blackbox]). In this work, we study the fairness properties of ML models subject to different attack types, with a particular focus on the realistic eyeglass attack.
@@
Different forms of defenses against adversarial examples have been proposed, including, but not limited to, defenses that detect attacks (e.g., [Metzen17Detector]); filter out adversarial perturbations (e.g., [Xu18Squeeze]); smoothen the classification boundaries to reduce model vulnerability (e.g., [Carlini23Certified, Cohen19RandSmooth, Salman19Robust]); verify robustness against specific adversaries (e.g., [Katz19Marabou]); and adversarially train models to inherently increase their robustness (e.g., [kurakin2017adversarial, madry2018towards]). Due to its intuitive nature, its ability to improve adversarial robustness in a practical manner against different attack types, and absence of impact on model’s inference time, adversarial training is particularly appealing and was widely studied (e.g., [goodfellow2015advtraining, Kurakin16AdvTrain, li2023wat, Lucas23AdvTrain, madry2018towards, Shafahi19PGD, tong2021facesec, Wong20Fast]). Accordingly, in this work, we focus on improving the fairness properties of adversarially trained models, especially against more realistic attacks. As adversarial training is often combined with other defenses (e.g., randomized smoothing [Salman19Robust]), we expect that our approach and findings would also apply to defenses other than adversarial training.
@@ 2.2 Fairness and Adversarial Robustness
Algorithmic fairness, in general, and fairness in ML, specifically, have been extensively studied in recent years [verma2018fairness]. Researchers have defined different fairness notions between individuals and groups, ranging from precision parity to recall parity [verma2018fairness], studied when notions can or cannot be satisfied simultaneously (e.g., [chouldechova2017fair]), and proposed means to promote different notions of fairness (e.g., [corbett2017algorithmic, singh2024representation]). In addition to ethical considerations that often motivate fairness, another key motivation behind fairness is improving the general performance of the system (e.g., benign accuracy), which sometimes stems as a side effect of improving fairness [karkkainenfairface].
Prior work has mainly focused on studying fairness in the context of benign inputs (e.g., [agarwal2018reductions, hashimoto2018fairness, karkkainenfairface]). However, a line of work has studied fairness in the context of adversarial robustness—particularly, disparities in adversarial robustness between different (source) classes and groups [nanda2021fairness, xu2021robust]. While defenses against adversarial examples such as adversarial training significantly improve robust accuracy, they do not guarantee fairness across classes or demographic groups. In high-stakes applications such as face recognition, this limitation is particularly concerning: a system may appear robust overall, yet still perform disproportionately worse on certain classes or subpopulations, exposing vulnerable groups to higher risks of misidentification or security breaches. In concurrent work, Nanda et al. [nanda2021fairness] and Xu et al. [xu2021robust] noticed that certain classes or groups may have significantly lower robust accuracy than others. These efforts highlighted the shortcomings of evaluating fairness only under benign conditions, and reinforced the need for defenses that explicitly address both robustness and fairness under attack. As noted by Rosenberg et al. [rosenberg2023fairness], similar disparities are also exhibited by face-recognition models where face-obfuscation with -norm bounded perturbations to attain privacy are less effective for minority groups.
@@
Several methods have emerged to tackle fairness in adversarial settings [jin2025enhancing, li2023wat, ma2022fat, sun2023improving, wei2023cfa, xu2021robust, zhang2024towards]. In particular, Fair Robust Learning (FRL) [xu2021robust] aims to improve worst-class robust accuracy through a min–max framework that reweights adversarial samples and adjusts attack strengths during training. Balance Adversarial Training (BAT) [sun2023improving] tackles fairness by encouraging uniformity in predicted classes and equalizing attack difficulty across source classes. Fairness-Aware Adversarial Learning (FAAL) [zhang2024towards] improves worst-class robustness by dynamically reweighting samples and using a form of so-called Distributionally Robust Optimization (DRO) to optimize worst-case performance. Last, Confusional Spectral Regularization (SpecNorm) [jin2025enhancing] penalizes the spectral norm of the confusion matrix to bound the worst-case class-level robust error. Similar to SpecNorm, Sy-FAR optimizes particular properties of the confusion matrix on adversarial examples (§4). However, Sy-FAR departs from norm-based constraints and instead focuses on directional imbalances in misclassification rates. SpecNorm and FAAL are considered the leading approaches for enhancing fair source-class adversarial robustness, thus we include them in our experiments (§5).
In contrast to prior work, Sy-FAR aims to improve fair adversarial robustness through novel means, by optimizing symmetry—while this property sometime naturally arises, misclassification patterns of adversarial examples between most class pairs are often asymmetric [mei2023towards]. Moreover, in this work we identify disparities in target-class adversarial robustness as a key limitation of existing techniques and show that Sy-FAR helps make it less severe. Last, previous work has mainly focused on predefined groups (e.g., genders or races), while we show how to improve fair adversarial robustness for any arbitrary group—addressing an issue that is known to be computationally challenging for certain notions of fairness.